Please help me solve this equation: 10x+3-20x=4-5x. I also need help on how to show my work on this problem. As well as on how to check this problem to see if this is the right solution.
Answer by stanbon(75887)

solve this equation:
10x+3-20x=4-5x.
-----
Combine the "x-terms" on the left side to get:
-10x+3 = 4-5x
---
Add 10x to both sides to get:
Subtract 4 from both sides to get:
5x = -1
Divide both sides by 5 to get:
----------
x = -1/5
------------------
Checking that answer in the original equation:
10x+3-20x=4-5x
-----------------------
10(-1/5)+3-(20)(-1/5) = 4 -5(-1/5)
-----------------------
-2 + 3 + 4 = 4 +1
-----------------------
5 = 5
==========================
x = -1/5 is the solution
===========================
Cheers,
Stan H.